# Adversarial Pruning: A Survey and Benchmark of Pruning Methods for Adversarial Robustness

### v1

@@ 4.1 Adversarial Pruning Evaluations
Table 2: AP evaluations. AA and PGD refer to AP evaluations using AutoAttack [13] and PGD [12], respectively. Additional attacks are reported under “Others", including the Fast Gradient Sign Method attack (FGSM) [46], Carlini&Wagner (CW) [63], Zeroth Order Optimization (ZOO) [64], DeepFool [65] and Brendel & Bethge (BB) [66]. In “Security curve" we report whether multiple perturbation sizes were used for the given attacks, while through “Iter$>$10" we report whether the PGD attacks were run with more than 10 iterations.
Name AA PGD Others Security Curve Iter$>$10
@@
In Table 2, we summarize the details of each AP adversarial evaluation and highlight their related issues. All in all, running adversarial attacks equals solving an optimization problem to find suitable perturbations that cause samples to be misclassified. Such as any optimization problem, an inaccurate hyperparameter optimization or setup (including the implementation) can lead to sub-optimal solutions, thus failing to provide a better estimate for the problem, as well known in the adversarial ML SoA [22, 23, 28]. It is, therefore, fundamental to assess the validity of the adversarial evaluations carried out in AP papers.
@@
AA, PGD and Other attacks. In columns “AA", “PGD", and “Others", we indicate whether the AP method evaluated the robustness using, respectively, the AutoAttack framework [13], PGD [12] or/and any other attack: through these columns, we directly highlight the diversity in the APs evaluations. AA, being an ensemble of four attacks, can be considered to supersede PGD: in fact, using different attacks at once can help ward off potential optimization problems, which are more likely using a single attack or few nearly identical versions [22] (e.g., PGD and FGSM). In addition, AA comprises a version of PGD (denominated APGD) that automatically optimizes the step size and uses two loss versions, regularly improving the results over standard PGD [12]. We label with AA every AP method testing at least some models using AutoAttack. We find a total of 9/26 APs using AA, while the remaining 17/26 use PGD as the predominant attack. [^1] Only a few AP methods were published before AA.
@@
Security curve. When using maximum-confidence attacks (e.g., PGD), the threat model encompasses a single perturbation budget $\epsilon$ (e.g., $\nicefrac{{8}}{{255}}$ with $\ell_{\infty}$ norm on the CIFAR10 dataset), thus returning the adversarial robustness of the model for that specific budget. However, evaluating the robustness of a single perturbation size limits the evaluation, as it is not clear how robustness decreases when increasing or decreasing the perturbation budget. Instead, a complete evaluation can be achieved through security curves, which plot the adversarial robustness against the perturbation norm [22, 27]. Using attacks such as PGD or AA, however, would require a substantial number of attack runs to plot such a curve, since a single attack run is bound to a fixed attack budget (see Sect. 2.2). In lieu of this demanding solution, minimum-norm attacks allow to straightforwardly plot the curve with a single attack run by returning adversarial examples not bounded to a single perturbation budget, but rather the smallest perturbation necessary [27]. Among the AP methods, although few also use minimum-norm attacks, none display security curves but rather clip the evaluation to a standard perturbation, with the unique exception of HYDRA, which plots the robustness over few discrete perturbation values with PGD attack [16], yet not using a minimum norm to display a full curve.
@@
Iter$>$10. When using single attacks such as PGD, it is crucial to carefully select the hyperparameters to ensure convergence [22]. Being PGD an iterative approach, although time-requiring, running the attack with a meaningful number of iterations improves the attack reliability [23]. We thus list in “Iter$>$10" the attacks using at least 10 iterations, which is a common default bare-minimum hyperparameter for the PGD attack, often leading to sub-optimal solutions. We find 3/26 setups using less than 10 iterations.
@@ 4.2 Benchmark Experimental Setting
Datasets and architectures. We focused on the CIFAR10 and SVHN datasets on two specific architectures, ResNet18 and VGG16. We found combinations between these networks and datasets to be the best compromise with the available implementations and the most common ones in the analyzed papers.
Sparsity rate. Given the difference in the effect of structured ($S$) vs unstructured ($US$) pruning, we selected two different sets of sparsity rates, considering the lower tolerance to high sparsities of structured implementations (which, therefore, imply choosing a lower range of $s_{r}$ values). Thus, when using $S$, we prune each model with 50%, 75% and 90& $s_{r}$, while when using $US$, we prune each model with 90%, 95% and 99% $s_{r}$.
Attacks. Following the discussion concerning the AP evaluations of Table 2, to re-evaluate the robustness of the AP methods on the fixed models, datasets and sparsity rates, we used the AutoAttack (AA) [13] ensemble. We run our evaluation on the entire test set, restricting to the $\ell_{\infty}$ norm threat model, and using $\nicefrac{{8}}{{255}}$ as perturbation budget $\epsilon$. To avoid having a single scalar robustness evaluation, we additionally used the Fast Minimum-Norm (FMN [27]) attack, through which we can collect multiple perturbation norms not confined to a single value and thus draw a complete security curve. In addition, to have a better estimate of the robustness throughout the entire perturbation norm scale, we optimized the hyperparameters of the FMN attack for each model under test following the HO-FMN procedure [28]. We show how the HO-FMN attack allows us to plot security curves, and we show an analysis of such curves on the CIFAR10 dataset and $US$ pruning, while more curves can be found in the available benchmark.
@@ 4.3 Contributing to the Benchmark
  • Describing the AP method pipeline and specifics in our dedicated Google Form. This will create a JSON entry with the novel AP description, which is automatically sent to the author.
@@
We will then evaluate the submission and update our leaderboard. Through our benchmark, we can evaluate the security curves of different AP methods for different sparsities and test every available checkpoint. Finally, in addition to novel AP submissions, we welcome existing checkpoints (from the AP authors whose results have been reproduced in this paper) when deemed not up to par with their AP potential.
@@ 4.4 Re-evaluation Results
Table 7: The pretrained models, evaluated with AutoAttack, for each of the A-methods. For PwoA [21], which distilles on a robust pretrained model, we used the HARP [20] robust checkpoints.
@@
Available implementations. We found a total of 11 available AP implementations, which we subdivided based on structured ($S$) and unstructured ($US$) pruning. Among these, we total 4 $S$ and 10 $US$ available pruning implementations, thus pruning 14 models for each of the 12 dataset/network/sparsity combinations for a total of 168 models. We point out that, although two further implementations were available, we could not solve their bugs or extend to the selected networks [57, 49]. Among the available ones, we specify that we experienced multiple issues in: (i) reproducing the results for MAD [38] and RST [53] on CIFAR10 with orders of 5 and 15 percentage points, respectively, and for which we welcome checkpoints used for the paper results; (ii) fitting the model for MAD [38] on SVHN, for which we had to modify the training procedures from the original one but failed in many occasions (hence the “-" entries in Table 5). In addition, SVHN was not originally implemented in multiple of the available APs; we thus extended the code in RADMM [17] (for which we occasionally had troubles in fitting the models), PwoA [21], RobustBird [54], FlyingBird [54] and Li20 [51]. In turn, we will mainly derive our conclusions and analysis from the CIFAR10 results, and then only validate on SVHN. To conclude, we specify that for HYDRA [16], leveraging additional data following [34], we used the standard dataset training to provide a fair comparison.
Pruning structure. On both the CIFAR10 and SVHN datasets, we notice from the results of Table 3 and Table 4, how $US$ pruning is less sensitive to higher sparsities than $S$. On the shared sparsity of 90%, except for TwinRep [56] on CIFAR10, all the AP methods hold a constantly greater accuracy and robustness against their structured counterpart. Pruning single weights as opposed to entire structures in fact, as widely known in the SoA [62], gives a higher flexibility and results in likewise higher performance.
@@
Is complexity rewarding? The analyzed AP methods are typically associated with complex pipelines and criterion designs. While comparing to a simpler pipeline is often demanding and, in general, not necessary, comparing to a simpler criterion is simpler and helps understand the true benefit of designing a complex and articulated pruning criterion. Therefore, considering LWM as a naive criterion for $US$ and filter L1 norm for $S$ [3, 61], we questioned how frequently papers adopting LIS, RELWM or SOLWM criteria compared with a naive one. In general, we found out to be quite rare for AP methods to compare to naive criteria [21, 57, 18, 42, 56, 20, 35, 59, 40, 43, 49]. We believe that such comparison helps validate the complexity of the AP and, most importantly, helps understand how much adopting such complex and often time-requiring criterion pays off.
@@
Flexible and cheap solution. We have shown how AP methods typically reach higher adversarial robustness when they use $G$ pruning, which is even higher when the $s_{r}$ in each layer is optimized. In addition, we have seen that most of the papers presenting AP methods do not compare their often complex criteria to a naive one, such as LWM. We thus question how much gain can be obtained by a naive and cheap criterion such as LWM when combined with a layer-wise strategy (i.e., different sparsity in each layer), and whether such gain could also be capable of surpassing more complex criterion using $L$ pruning. Therefore, we combined LWM with a different $s_{r}$ strategy for each layer. However, instead of optimizing the layerwise $s_{r}$ such as in HARP (which would clearly increase performances but also complexity), we make use of cheaper approaches such as Layer-wise Adaptive Magnitude Pruning LAMP [67], which computes an ideal layerwise sparsity without any additional cost, thus preserving the low-cost of a naive criterion but improving the performances of LWM. Interestingly, through our results in Table 8, we show that LWM has a significant gain from using the $G$ strategy of LAMP, to the extent that some of the more complex and time-requiring APs retain lower robustness. This corroborates results from prior work (HARP and Heracles [35, 20]) showing the benefits of non-uniformity on their corresponding competing AP methods [35, 20]. Most importantly, it shows that complex AP methods can often be surpassed by simple and low complexity solutions increasing the “flexibility" of the chosen weights to be pruned.
@@
Unreliable evaluations. While many of the published AP methods do not test robustness with AA, we find only two methods among the available implementations never testing any model on AA, RADMM [17] and Li20 [51]. However, for the rest of the available APs testing on AA, we notice how not every paper extensively tests all model-sparsity combinations on AA, but rather executes just a few trials on selected pairs of models while only relying on PGD in an extensive way. Therefore, by extending the evaluations using AA, it is possible to show how APs have often overestimated robustness. Still, it is hard to exactly compare the results of our re-evaluation with the original papers’ results, since models in their papers are pruned to sparsities which do not necessarily match our benchmark setting. We specify that our benchmark works right toward this direction: providing a template for previous and new APs to be fairly and reliably tested and compared. Despite these limitations, we report results from the original AP papers for similar (or the same, when possible) sparsities on CIFAR10, aiming to quantify the extent to which robustness has been overestimated in available APs.
@@
Table 9: Summary of robustness overestimation in AP papers for CIFAR10 for both $US$ and $S$ pruning. We consider, for each AP, the model with $s_{r}$ matching ours, or when not available, the closest and lower of our sparsities (i.e., $93.75$ is compared to our $90$). We indicate with “Orig." the paper’s PGD robustness, with “Ours" the one computed with AA, and the difference between the two.
@@
Name $s_{r}$ Orig. Ours Diff. Name $s_{r}$ Orig. Ours Diff.
@@
Security evaluation curve. To have a comprehensive security evaluation on multiple perturbation norms and thus plot security curves, we use the SoA minimum-norm attack approach from HO-FMN [28], which optimizes the attack hyperparameters to find the best configuration on which to run the attack for the model under test. Using a minimum-norm attack, the robustness evaluation is not bounded to a single scalar value (e.g., $\nicefrac{{8}}{{255}}$), since the goal is to find the smallest perturbation norm. Therefore, as discussed in Sect. 2.2, the adversarial examples found by the attack are associated with multiple perturbations, which makes the evaluation of a security curve straightforward and cheap [22, 27].
Figure 2: The security curves of HARP, HYDRA, and FlyingBird for each sparsity rate with $US$ pruning on CIFAR10. The plots have been created using the HO-FMN optimization procedure for the FMN attack [27, 28]. The curves show robust accuracy against the perturbation norm $\epsilon$, where the value $\nicefrac{{8}}{{255}}$ is represented with a vertical dotted line. The bar subplot shows the colour of the most robust model, describing how the “leader model" varies over perturbation.
In reference to HO-FMN, we run the search for the attack hyperparameters on 2048 samples for each pruned model. We then run the attack on the remaining samples using the best hyperparameters found on the DLR loss, SGD optimizer and Cosine Annealing scheduler. We show, in Figure 2, the security curves for the top-3 APs on the CIFAR10 dataset, HYDRA, HARP and FlyingBird, for each of the three sparsities of $US$. Security curves show the decrease of robustness as the perturbation grows, which does not necessarily reflect the robustness evaluation computed on a scalar perturbation (e.g., through AA). To highlight such an aspect, we show a bar subplot in Figure 2 using the colour of the most robust model, thus showing how, as the perturbation grows, the robustness across models can change significantly. When the perturbation is small, we find HARP to be the most robust model. Instead, FlyingBird is typically the most robust around just a small window around $\nicefrac{{8}}{{255}}$, which is the typical perturbation norm used to evaluate the models. As the perturbation grows, HYDRA is instead constantly the most robust model. Using security curves represents a thus fundamental tool to guarantee a complete adversarial robustness evaluation and shows how a single scalar evaluation might not always be sufficient.

### v2

@@ 4.1 Adversarial Pruning Evaluations
Table 2: AP evaluations. AA and PGD refer to AP evaluations using AutoAttack [13] and PGD [12], respectively. Additional attacks are reported under “Others”, including the Fast Gradient Sign Method attack (FGSM), Carlini&Wagner (CW), Zeroth Order Optimization (ZOO), DeepFool and Brendel & Bethge (BB) [63]. In “Robustness curve” we report whether multiple perturbation sizes were used for the given attacks, while through “Iter$>$10” we report whether the PGD attacks were run with more than 10 iterations.
Name AA PGD Others Robustness Curve Iter$>$10
@@
In Table 2, we summarize the details of each AP adversarial evaluation and highlight their related issues. All in all, running adversarial attacks equals solving an optimization problem to find suitable perturbations that cause samples to be misclassified. Such as any optimization problem, an inaccurate hyperparameter optimization or setup (including the implementation) can lead to sub-optimal solutions, thus failing to provide a better estimate for the problem, as well known in the adversarial ML literature [25, 62, 26]. It is, therefore, fundamental to assess the validity of the adversarial evaluations carried out in AP papers.
@@
AA, PGD and Other attacks. In columns “AA”, “PGD”, and “Others”, we indicate whether the AP method evaluated the robustness using, respectively, the AutoAttack framework [13], PGD [12] or/and any other attack. Through these columns, we directly highlight the diversity in the APs evaluations. AA, being an ensemble of four attacks, can be considered to supersede PGD; indeed, using multiple attacks at once can help ward off potential optimization issues, which are more likely using a single attack or few nearly identical versions [25]. In addition, AA comprises a version of PGD (named AutoPGD) that automatically optimizes the step size and uses two loss versions, typically outperforming standard PGD [12]. We label with AA every AP method tested with AutoAttack. We find a total of 9/26 AP methods tested using AA, while the remaining 17/26 use PGD as the predominant attack. [^2] Note, however, that only a few AP methods were published before AA.
@@
Robustness curve. When using maximum-confidence attacks (e.g., PGD), the threat model encompasses a single perturbation budget $\epsilon$ (e.g., $\nicefrac{{8}}{{255}}$ with $\ell_{\infty}$ norm on the CIFAR10 dataset), thus returning the adversarial robustness of the model for that specific budget. However, evaluating adversarial robustness at a single perturbation size limits the evaluation, as it is not clear if robustness decreases more or less gracefully when increasing the perturbation size. Instead, a complete evaluation can be achieved through robustness curves, which plot the adversarial robustness against the perturbation norm [15, 25, 24]. Using attacks such as PGD or AA, however, would require a substantial number of attack runs to plot such a curve, since a single attack run is bound to a fixed attack budget (see Sect. 2.2). In lieu of this demanding solution, minimum-norm attacks allow to straightforwardly plot the curve with a single attack run by returning adversarial examples not bounded to a single perturbation budget, but rather the smallest perturbation necessary [24]. Among the AP methods, although few also use minimum-norm attacks, none display robustness curves but rather clip the evaluation to a standard perturbation, with the unique exception of HYDRA, which plots the robustness over few discrete perturbation values with PGD attack [16], yet not using a minimum norm to display a full curve.
@@
Iter$>$10. When using single attacks such as PGD, it is crucial to carefully select the hyperparameters to ensure convergence [25]. Being PGD an iterative approach, although time-requiring, running the attack with a meaningful number of iterations improves the attack reliability [62]. We thus list in “Iter$>$10” the attacks using at least 10 iterations, which is a common default bare-minimum hyperparameter for the PGD attack, often leading to sub-optimal solutions. We find 3/26 setups using less than 10 iterations.
@@ 4.2 Benchmark Experimental Setting
Datasets and architectures. We mainly focus on the CIFAR10 and SVHN datasets on two specific architectures, ResNet18 and VGG16. We found combinations between these networks and datasets to be the best compromise with the available implementations and the most common ones in the analyzed papers.
Sparsity rate. Given the difference in the effect of structured ($S$) vs unstructured ($US$) pruning, we selected two different sets of sparsity rates, considering the lower tolerance to high sparsities of structured implementations (which, therefore, implies choosing a lower range of $s_{r}$ values). Thus, when using $S$, we prune each model with 50%, 75%, and 90& $s_{r}$, while when using $US$, we prune each model with 90%, 95%, and 99% $s_{r}$.
Attacks. Following the discussion concerning the AP evaluations of Table 2, to re-evaluate the robustness of the AP methods on the fixed models, datasets, and sparsity rates, we used the AutoAttack (AA) ensemble [13]. We run our evaluation on the entire test set, restricting to the $\ell_{\infty}$ norm threat model, and using $\nicefrac{{8}}{{255}}$ as perturbation budget $\epsilon$. To avoid having a single scalar robustness evaluation, we additionally used the Fast Minimum-Norm (FMN) attack [24], through which we can collect multiple perturbation norms not confined to a single value and thus draw a complete robustness curve. In addition, to have a more reliable estimate of the robustness, we optimized the hyperparameters of FMN for each model under test following the HO-FMN procedure [26]. We show how HO-FMN allows us to plot robustness curves, and report an analysis of such curves on the CIFAR10 dataset and $US$ pruning, while more curves can be found in the available benchmark.
@@ 4.3 Contributing to the Benchmark
  • 1.
    Describing the AP method pipeline and specifics in the dedicated section, which will generate a JSON file.
  • 2.
@@
(i) We will then evaluate the submission and update our leaderboard. Through our benchmark, we can evaluate the robustness curves of different AP methods for different sparsities and test every available checkpoint. Finally, in addition to novel AP submissions, we welcome existing checkpoints (from the AP authors whose results have been reproduced in this paper) when deemed not up to par with their AP potential.
@@ 4.4 Re-evaluation Results
Table 7: The pretrained models, evaluated with AutoAttack, for each of the A-methods. For PwoA [21], which distills on a robust pretrained model, we used the HARP [20] robust checkpoints.
@@
Available implementations. We found a total of 11 available AP implementations, which we subdivided based on structured ($S$) and unstructured ($US$) pruning. Among these, we total 4 $S$ and 10 $US$ available pruning implementations, thus pruning 14 models for each of the 12 dataset/network/sparsity combinations for a total of 168 models. We point out that, although two further implementations were available, we could not solve their bugs or extend to the selected networks [56, 48]. Among the available ones, we specify that we experienced multiple issues in: (i) reproducing the results for MAD [37] and RST [52] on CIFAR10 with orders of 5 and 15 percentage points, respectively, and for which we welcome checkpoints used for the paper results; (ii) fitting the model for MAD [37] on SVHN, for which we had to modify the training procedures from the original one but failed in many occasions (hence the “-” entries in Table 5). In addition, SVHN was not originally implemented in multiple of the available APs; we thus extended the code in RADMM [17] (for which we occasionally had troubles in fitting the models), PwoA [21], RobustBird [53], FlyingBird [53] and Li20 [50]. In turn, we will mainly derive our conclusions and analysis from the CIFAR10 results and then only validate on SVHN. To conclude, we specify that for HYDRA [16], leveraging additional data following [33], we used the standard dataset training to provide a fair comparison.
Pruning structure. On both CIFAR10 and SVHN, we notice from the results of Table 3 and Table 4, how $US$ pruning is less sensitive to higher sparsities than $S$. On the shared sparsity of 90%, except for TwinRep [55] on CIFAR10, all the AP methods hold a constantly greater accuracy and robustness against their structured counterpart. Pruning single weights as opposed to entire structures, as widely known in the field [61], gives indeed higher flexibility and results in likewise higher performance.
@@
Is complexity rewarding? The analyzed AP methods are typically associated with complex pipelines and criterion designs. While comparing to a simpler pipeline is often demanding and, in general, not necessary, comparing to a simpler criterion is simpler and helps understand the true benefit of designing a complex and articulated pruning criterion. Therefore, considering LWM as a naive criterion for $US$ and filter L1 norm for $S$ [3, 60], we questioned how frequently papers adopting LIS, RELWM, or SOLWM criteria compared with a naive one. In general, we found out to be quite rare for the surveyed AP methods to compare to naive criteria. We believe that such comparison helps validate the complexity of the AP and, most importantly, helps understand how much adopting such complex and often time-requiring criterion pays off.
@@
Flexible and cheap solution. We have shown how AP methods typically reach higher adversarial robustness when they use $G$ pruning, which is even higher when the $s_{r}$ in each layer is optimized. In addition, we have seen that most of the papers presenting AP methods do not compare their often complex criteria to a naive one, such as LWM. We thus question how much gain can be obtained by a naive and cheap criterion such as LWM when combined with a layer-wise strategy (i.e., different sparsity in each layer), and whether such gain could also be capable of surpassing more complex criterion using $L$ pruning. Therefore, we combined LWM with a different $s_{r}$ strategy for each layer. However, instead of optimizing the layerwise $s_{r}$ such as in HARP (which would clearly increase performances but also complexity), we make use of cheaper approaches such as Layer-wise Adaptive Magnitude Pruning LAMP [64], which computes an ideal layerwise sparsity without any additional cost, thus preserving the low-cost of a naive criterion but improving the performances of LWM. Interestingly, through our results in Table 8, we show that LWM has a significant gain from using the $G$ strategy of LAMP, to the extent that some of the more complex and time-requiring APs retain lower robustness. This corroborates results from prior work (HARP and Heracles [34, 20]) showing the benefits of non-uniformity on their corresponding competing AP methods [34, 20]. Most importantly, it shows that complex AP methods can often be surpassed by simple and low complexity solutions increasing the “flexibility” of the chosen weights to be pruned.
@@
Unreliable evaluations. While many of the published AP methods do not test robustness with AA, we find only two methods among the available implementations never testing any model on AA, RADMM [17] and Li20 [50]. However, for the rest of the available APs testing on AA, we notice how not every paper extensively tests all model-sparsity combinations on AA, but rather executes just a few trials on selected pairs of models while only relying on PGD in an extensive way. Therefore, by extending the evaluations using AA, it is possible to show how APs have often overestimated robustness. Still, it is hard to exactly compare the results of our re-evaluation with the original papers’ results, since models in their papers are pruned to sparsities that do not necessarily match our benchmark setting. We specify that our benchmark works right toward this direction: providing a template for previous and new APs to be fairly and reliably tested and compared.
Despite these limitations, we report results from the original AP papers for similar (or the same, when possible) sparsities on CIFAR10, aiming to quantify the extent to which robustness has been overestimated in available APs.
@@
Table 9: Summary of robustness overestimation in AP papers for CIFAR10 for both $US$ and $S$ pruning. We consider, for each AP, the model with $s_{r}$ matching ours, or when not available, the closest and lower of our sparsities (i.e., $93.75$ is compared to our $90$). We indicate the paper’s PGD robustness with “Orig.” and with “Ours,” the one computed with AA, and the difference between the two.
@@
Name $s_{r}$ Orig. Ours Diff. Name $s_{r}$ Original Ours Diff.
@@
Robustness curves. To have a comprehensive evaluation on multiple perturbation norms and thus plot robustness curves, we use one of the most recent minimum-norm attacks, known as HO-FMN [26], which optimizes the attack hyperparameters to find the best configuration on which to run the attack for the model under test. Using a minimum-norm attack, the robustness evaluation is not bounded to a single scalar value (e.g., $\nicefrac{{8}}{{255}}$), since the goal is to find the smallest perturbation norm. Therefore, as discussed in Sect. 2.2, the adversarial examples found by the attack are associated with multiple perturbations, which makes the evaluation of the curve straightforward and efficient [25, 24].
Figure 3: The robustness curves of HARP, HYDRA, and FlyingBird for each sparsity rate with $US$ pruning on CIFAR10. The plots have been created using the HO-FMN optimization procedure for the FMN attack [24, 26]. The curves show robust accuracy against the perturbation norm $\epsilon$, where the value $\nicefrac{{8}}{{255}}$ is represented with a vertical dotted line. The bar subplot shows the color of the most robust model, describing how the “leader model” varies over perturbation.
In reference to HO-FMN, we run the search for the attack hyperparameters on 2048 samples for each pruned model. We then run the attack on the remaining samples using the best hyperparameters found on the DLR loss, SGD optimizer and Cosine Annealing scheduler. We show, in Figure 3, the robustness curves for the top-3 APs on the CIFAR10 dataset, HYDRA, HARP, and FlyingBird, for each of the three sparsities of $US$. Robustness curves show the decrease of robustness as the perturbation grows, which does not necessarily reflect the robustness evaluation computed on a scalar perturbation (e.g., through AA). To highlight such an aspect, we show a bar subplot in Figure 3 using the color of the most robust model, thus showing how, as the perturbation grows, the robustness across models can change significantly. When the perturbation is small, we find HARP to be the most robust model. Instead, FlyingBird is typically the most robust around just a small window around $\nicefrac{{8}}{{255}}$, which is the typical perturbation norm used to evaluate the models. As the perturbation grows, HYDRA is instead constantly the most robust model. Using robustness curves represents a thus fundamental tool to guarantee a complete adversarial robustness evaluation and shows how a single scalar evaluation might not always be sufficient.
Extension on ImageNet dataset. We additionally expand on the ImageNet dataset and ResNet50 architecture for the HARP [20] and HYDRA [16] methods, resulting in a comparison of AP methods for a bigger-scale dataset. As for the other datasets, we allow the online benchmark to host ImageNet models, thus providing a further comparison for AP methods. We show, in Table 10, the results for $US$ pruning, which, similar to the results for different datasets and architectures, confirm the benefits of the optimized layer-wise sparsity of HARP, compared to HYDRA.
Table 10: The re-evaluation results for $US$ pruning methods on Imagenet and ResNet50.
        Sparsity         HARP         HYDRA
        90%         23.35         21.36
        95%         22.88         19.71
        99%         12.56         10.37